# Glauber predictions for oxygen and neon collisions at the LHC

### v2

@@ Abstract
The Glauber model is a widely used framework for describing the initial conditions in high-energy nuclear collisions. TGlauberMC is a Monte Carlo implementation of this model that enables detailed, event-by-event calculations across various collision systems. In this work, we present an updated version of TGlauberMC (v3.3), which incorporates recent theoretical developments and improved parameterizations, especially relevant for small collision systems. We focus on the oxygen–oxygen (OO), neon–neon (NeNe), and proton-oxygen (pO) collisions at the Large Hadron Collider (LHC) in July 2025, where precise modeling of nuclear geometry and fluctuations is essential. The updated version includes revised nuclear density profiles and an enhanced treatment of nucleon substructure. We calculate geometrical cross sections for all relevant collision systems and explore initial-state observables, providing predictions for particle production trends at $\sqrt{s_{\rm NN}}=5.36$ TeV. In particular, we present a prediction for the centrality dependence of mid-rapidity multiplicity in OO and NeNe collisions. The updated code is publicly available to support the heavy-ion community with a robust and flexible tool for studying strongly interacting matter in small and intermediate-sized nuclear systems.
@@ I Introduction
The study of heavy-ion collisions at ultra-relativistic energies has been instrumental in exploring the properties of the quark-gluon plasma (QGP), a state of matter believed to have existed shortly after the Big Bang Busza:2018rrf . Experiments at both RHIC and the LHC have provided compelling evidence for the formation of QGP in large collision systems such as AuAu and PbPb at high collision energies BRAHMS:2004adc ; PHENIX:2004vcz ; PHOBOS:2004zne ; STAR:2005gfr ; ALICE:2022wpn ; CMS:2024krd . Among the most prominent observables signaling QGP formation are collective flow phenomena, characterized by anisotropic flow coefficients such as $v_{2}$ and higher moments Heinz:2013th , and the phenomenon of jet quenching, in which high-energy partons lose energy as they traverse the dense medium Cunqueiro:2021wls . Surprisingly, several QGP-like features have also been observed in small collision systems, such as proton-lead (pPb) and even high-multiplicity proton-proton (pp) collisions Loizides:2016tew ; Nagle:2018nvi ; Grosse-Oetringhaus:2024bwr . These include long-range correlations and non-zero elliptic flow coefficients, raising questions about the minimal conditions required for QGP formation. However, one signature remains notably absent in small systems: jet quenching. Despite observing various collective phenomena, experiments have not identified energy loss of high-$p_{T}$ partons in pPb collisions ALICE:2017svf ; ATLAS:2022iyq . To further elucidate the system-size dependence of QGP signatures and to search for the possible onset of jet quenching in intermediate-sized systems, oxygen-oxygen (OO) and neon-neon (NeNe) collisions at a center-of-mass energy ($\sqrt{s_{\rm NN}}$) of 5.36 TeV per nucleon pair provided by the LHC in July, 2025 are of great interest Citron:2018lsq ; Brewer:2021kiv . These collision systems provide a crucial bridge between small and large systems, allowing precise control measurements of observables such as jet quenching and elliptic flow Lim:2018huo ; Sievert:2019zjr ; Huang:2019tgz ; Huss:2020dwe ; Jia:2022ozr . Specifically, the role of $\alpha$ clusters in small nuclei including 16O has long been discussed and may be imprinted in the final-state flow measurements Rybczynski:2017nrx ; Rybczynski:2019adt ; Li:2020vrg ; Summerfield:2021oex ; Ding:2023ibq ; YuanyuanWang:2024sgp ; Zhang:2024vkh . Recently, comparing $v_{2}$ in NeNe to $v_{2}$ in OO collisions was highlighted to provide a precision test of the role of initial geometry and the QGP paradigm Giacalone:2024luz . Additionally, proton–oxygen (pO) collisions at a $\sqrt{s_{\rm NN}}=9.62$ TeV were provided, primarily to improve hadronic interaction models used in the interpretation of extensive air showers in cosmic-ray physicsDembinski:2020dam , but they also present a unique opportunity to investigate QGP-related phenomena.
@@
Interpretation of high-energy nucleus collision data typically depends on the use of classical Glauber Monte Carlo (MCG) models to describe the initial condition Miller:2007ri ; dEnterria:2020dwq . The general approach is to deduce properties of QCD matter by measuring the distributions of a variety of observables in AA collisions, and comparing them to the same measurements in proton-proton (pp) collisions, where a QGP is not expected to be produced. To compare colliding systems of different sizes requires hence appropriate normalization of the measured distributions by using e.g. the number of participating nucleons ($N_{\rm part}$), the number of independent binary nucleon-nucleon (NN) collisions ($N_{\rm coll}$), the medium transverse area ($A_{\perp}$), or the overlap eccentricity $\epsilon_{n}$ (given by the $n^{\rm th}$ moments of its azimuthal spatial distribution).
@@
Widely used and publicly available MCG models are the TGlauberMC Loizides:2014vua ; Loizides:2016djv ; Loizides:2017ack , which developed from the original PHOBOS implementation Alver:2008aq and GLISSANDO codes Broniowski:2007nz ; Rybczynski:2013yba ; Bozek:2019wyr . The Glauber initial condition is widely used in simulations of heavy-ion collisions as provided by TRENTO Moreland:2014oya , Trajectum Nijs:2021clz and MUSIC Paquet:2015lta , as well as by event generators such as HIJING Wang:1991hta , AMPT Lin:2004en , URQMD Bass:1998ca and EPOS Werner:2010aa . An alternative description of the initial condition is provided by IP-Glasma Schenke:2018fci .
@@
In this article, we use an updated version (v3.3) of the TGlauberMC model to predict properties of OO, NeNe, as well as pPb and PbPb collisions at 5.36 TeV, and pO collisions at 9.62 TeV. In particular, we focus on the calculation of the expected cross sections and discuss aforementioned associated initial-state quantities, such as $N_{\rm part}$, $N_{\rm coll}$ and eccentricity distributions for OO and NeNe collisions including the ratio of eccentricities for NeNe to OO as a proxy for the respective ratio of $v_{2}$ in these two systems. Finally, we also discuss particle production at 5.36 TeV by outlining some features derived from the multiple parton interaction (MPI) picture Sjostrand:2017cdm , and use existing data at 5.36 TeV to predict the multiplicity in OO and NeNe collisions. As for previous versions of the model, the source code for TGlauberMC (v3.3) has been made publicly available at HepForge glaucode .
@@ II The Monte Carlo Glauber approach
Monte Carlo Glauber (MCG) calculations evaluate the cross section of two nuclei with $A$ and $B$ nucleons stochastically by distributing them in coordinate space according to the corresponding nuclear densities, separated by an impact parameter $b$ sampled from $\mathrm{d}\sigma/\mathrm{d}b\propto b$. For simplicity, the impact-parameter variable $b$ or $b_{\rm NN}$ denotes the distance between nuclei (or nucleons) in the transverse plane. The nuclear transverse profiles of the nuclei are usually approximated with parametrizations of charge density distributions extracted from low-energy electron-nucleus scattering experiments DeVries:1987atn . For large nuclei, they are usually described by a 3-parameter Fermi (3pF) distribution
@@
with half-density radius $R$ and diffusivity $a$, and $w\neq 0$ describing small deviations from a spherical shape. Following the eikonal approximation, the collision is treated as a sequence of independent binary nucleon-nucleon collisions, i.e. the nucleons travel on straight-line trajectories and their interaction probability does not depend on the number of collisions they have suffered before. In the so-called “hard-sphere” (HS) approximation, an interaction takes place between two nucleons if the transverse distance $d$ (or generally the transverse nucleon–nucleon impact parameter $b_{\rm NN}$) between their centers satisfies $d<\sqrt{\sigma_{\rm NN}/\pi}$.
In the MCG approach, the number of participants and collisions are computed by counting the number of nucleons with at least one collision, and the total number of nucleon–nucleon collisions, respectively. The collision cross section is the fraction of simulated collisions with $N_{\rm coll}>0$ of all simulated collisions. Quantities such as the eccentricity PHOBOS:2006dbo , the triangularity Alver:2010gr , and higher moments Teaney:2010vd of the collision region at impact parameter $b$ are given by
@@
where $n$ denotes the moment, $r=\sqrt{x^{2}+y^{2}}$ and $\psi=\tan^{-1}\frac{y}{x}$. The averages are computed by considering the central positions of either, participant nucleons or binary nucleon–nucleon collisions, or of an admixture of the two, and additionally may be smeared to model matter production PHOBOS:2007vdf .
@@ III Transverse overlap function
Several extensions to the hard-sphere approximation have been explored to account for the fact that the nucleon matter profile is not a black disc dEnterria:2010xip . A commonly applied extension, referred to as TRENTO Moreland:2014oya , is to assume the nuclear matter profile of a nucleon to be a Gaussian with a certain width $w$, so that the nuclear overlap function
@@
is also Gaussian, i.e. $m=2$ and $B=2w$. For some settings (“bProfile = 3” Sjostrand:2017cdm ) PYTHIA assumes $B=1$ fm as characteristic size of the proton, and uses $m$ to interpolate between Exponential ($m\to 1$) and Gaussian ($m\to 2$) shape. The widely-used “Monash” tune sets $m=1.85$ Skands:2014pea . HIJING Wang:1991hta uses an overlap function based on modified Bessel functions of the 3rd kind,
@@
with $\mu={1.5}/{\sqrt{\sigma_{\rm NN}}}$. In the Eikonal formalism, the probability to interact is then given as $p(b_{\rm NN})=1-\exp\left(-k\,T_{\rm NN}(b_{\rm NN})\right)$ where $k$ is obtained by requiring $2\pi\,\int p(b_{\rm NN})\,b_{\rm NN}\,{\rm d}b_{\rm NN}=\sigma_{\rm NN}$. At LHC collision energies, a practical parametrization Rybczynski:2013mla based on the Euler $\Gamma(z)$ and incomplete $\Gamma(\alpha,z)$ functions, in the following also called $\Gamma$-parametrization, is
@@
Figure 2: Nuclear charge density for 16O versus radius for various descriptions compared to data from electron scattering Sick:1970ma . As explained in the text, “Oho2” is a parameterization based on the HO model, whose parameters are fit to the data, while “Odat” was already provided in Sick:1970ma . “Osat” is the exact result of the NNLOsat calculation Soma:2019bso ), and “Opar2” the respective 3pF fit. “Opar” and “Oho” profiles were previously included in TGlauberMC. The distributions are normalized as $\int\rho\,r^{2}\,{\rm d}r=Ze$. The respective $r_{\rm rms}$ values are given in the legend, denoted as $s$.
@@ IV Nuclear density profiles for oxygen and neon
The standard method to construct the nucleon configurations for MCG calculations is to draw the nucleon positions from $r^{2}\rho$, where the nuclear density $\rho$ may include angular deformations Loizides:2014vua . To mimic a hard-core repulsion potential between nucleons, a minimum inter-nucleon separation ($d_{\rm min}$) of $0.4$ fm between their centers is enforced when sampling the positions of the nucleons inside a nucleus. In order to ensure that the center-of-mass of each constructed nucleus is at $(0,0,0)$, the nucleons are individually “shifted” by the necessary amount, after all nucleon positions were sampled. Both, imposing the hard-core repulsion as well as the re-centering, induce a bias from the original $\rho$ distribution. As discussed in Ref. Loizides:2017ack , this can be overcome by a “reweighting” procedure, modifying the original $\rho$ distribution to counteract the bias. However, for $d_{\rm min}=0.4$ fm, and recentering via rotation around the $z$-axis (recentering option 4 in TGlauberMC), residual differences are negligible.
Table 1: Nuclear parameters for oxygen and neon parameterizations provided by TGlauberMC (v3.3) with given names as in the code. The 3pF parameterization is given in Eq. (1). The HO parameterization is given in Eq. (6). The last two configurations used deformed profiles, with deformation values as given in the comment field, and defined in Trajectum Giacalone:2024luz . The prefix “TR” indicates the profile is taken from Trajectum, where the name following the prefix denotes the system and version number as used in Trajectum. Included as “Osat” is also the exact profile from the ${\rm NNLO}_{\rm sat}$ chiral Hamiltonian calculation Soma:2019bso . For completeness, “O” is also listed which uses explicit nucleon configurations Lonardoni:2017hgs as introduced in Lim:2018huo .
@@
Odat 16 8 2.608 0.513 -0.051 Original parametrization (3pF plus corrections) from Sick:1970ma
Opar 16 8 2.608 0.513 -0.051 3pF parameterization DeVries:1987atn (same as TR_OXYGEN_V3)
Opar2 16 8 1.850 0.497 0.912 3pF parameterization (TR_OXYGEN_V4, fit to ${\rm NNLO}_{\rm sat}$ Soma:2019bso )
Oho 16 8 1.544 1.833 - Harmonic Oscillator DeVries:1987atn
Oho2 16 8 1.506 1.819 - Harmonic Oscillator (new fit to data from Sick:1970ma )
Osat 16 8 - - - Exact NNLOsat profile from Soma:2019bso , use instead of Opar2
O 16 8 - - - Nucleon configurations Lim:2018huo  (same as in TGlauberMC v2.7)
Ne 20 10 2.805 0.571 - 2pF parametrization DeVries:1987atn
Ne2 20 10 2.740 0.572 - 2pF parametrization DeVries:1987atn
Ne3 20 10 2.791 0.698 -0.168 3pF parametrization DeVries:1987atn
NeTr2 20 10 2.805 0.571 - TR_NEON_V2, (same as Ne with $\beta_{2}=0.721$ from Pritychenko:2013gwa )
@@
Figure 2 shows the nuclear charge density for 16O for data from electron scattering Sick:1970ma , including the original parameterization (“Odat”), as well as 3pF distributions with various parameters available in the literature (“Opar”, and “Opar2”). Ref. DeVries:1987atn also gives a parameterization based on the (modified) Harmonic Oscillator (HO) model
@@
labeled “Oho” in the figure. For this work, we also provide a new HO fit to the data, called “Oho2”, leading to $\alpha=1.506\pm 0.023$ and $1.819\pm 0.004$ with $\chi^{2}/n_{\rm dof}\approx 1$ assuming about 1% relative error on the data points. While it overall has a worse $\chi^{2}$ than “Odat”, it only has two parameters and can be used as alternative and allows systematic variation using the uncertainties provided on the fit parameters. To reflect the charge density, the distributions shown in Fig. 2 are normalized as $\int\rho\,r^{2}\,{\rm d}r=Ze$. A detailed list of all parameterizations shown in Fig. 2 is given in Tab. 1. The $r_{\rm rms}=\sqrt{\langle r\rangle}$ values are reported in the legend of the figure, and differ by nearly up to 5%. The reason is that the data do not constrain the tail of the distribution, which was found to affect the $r_{\rm rms}$ by a few percent, and thus experimental results between $2.65\pm 0.04$ and $2.73\pm 0.03$ have been reported Sick:1970ma . Furthermore small differences can also be expected from inclusion of neutrons in the density distributions, and the use of point-like protons ignoring the proton form factor. However, the distribution “Opar2” which results from a fit to the ${\rm NNLO}_{\rm sat}$ calculation Soma:2019bso exhibits a significantly larger $r_{rms}$ than the data (and than reported in Soma:2019bso ), and hence is not further considered here. Instead, the exact nuclear density obtained from the ${\rm NNLO}_{\rm sat}$ chiral Hamiltonian calculation Soma:2019bso is included and also shown in the figure. The nuclear charge density directly calculated by ${\rm NNLO}_{\rm sat}$ perfectly agrees with the data (also shown in the figure).
Table 2: Oxygen-16 and Neon-20 NLEFT and PGCM calculations taken from Trajectum Giacalone:2024luz .
@@
TR_OXYGEN_V15 PGCM, constraints imposed on projected states, with explicit clustering

### v3

# Glauber predictions for oxygen and neon collisions at energies available at the LHC

@@ Abstract
The Glauber model is a widely used framework for describing the initial conditions in high-energy nuclear collisions. TGlauberMC is a Monte Carlo implementation of this model that enables detailed, event-by-event calculations across various collision systems. In this work, I present an updated version of TGlauberMC (v3.3), which incorporates recent theoretical developments and improved parameterizations, especially relevant for small collision systems. I focus on the oxygen–oxygen (OO), neon–neon (NeNe), and proton-oxygen (pO) collisions at the Large Hadron Collider (LHC) in July 2025, where precise modeling of nuclear geometry and fluctuations is essential. The updated version includes revised nuclear density profiles and an enhanced treatment of nucleon substructure. Geometrical cross sections for all relevant collision systems are calculated and initial-state observables are explored to provide predictions for particle production trends at $\sqrt{s_{\rm NN}}=5.36$ TeV. In particular, a prediction for the centrality dependence of mid-rapidity multiplicity in OO and NeNe collisions is obtained. The updated code is publicly available to support the heavy-ion community with a robust and flexible tool for studying strongly interacting matter in small and intermediate-sized nuclear systems.
@@
v3, 03 Feb 2026: published version, code v3.3.2 includes smearing with TRENTO, and static functions
@@ I Introduction
The study of heavy-ion collisions at ultra-relativistic energies has been instrumental in exploring the properties of the quark-gluon plasma (QGP), a state of matter believed to have existed shortly after the Big Bang Busza et al. (2018). Experiments at the Relativistic Heavy Ion Collider (RHIC) and the LHC have provided compelling evidence for the formation of QGP in large collision systems such as gold–gold (AuAu) and lead–lead (PbPb) at high collision energies Arsene and others (2005); Adcox and others (2005); Back and others (2005); Adams and others (2005); Acharya and others (2024b); Hayrapetyan and others (2025c). Among the most prominent observables signaling QGP formation are collective flow phenomena, characterized by anisotropic flow coefficients such as $v_{2}$ and higher moments Heinz and Snellings (2013), and the phenomenon of jet quenching, in which high-energy partons lose energy as they traverse the dense medium Cunqueiro and Sickles (2022). Surprisingly, several QGP-like features have also been observed in small collision systems, such as proton-lead (pPb) and even high-multiplicity proton-proton (pp) collisions Loizides (2016a); Nagle and Zajc (2018); Grosse-Oetringhaus and Wiedemann (2024). These include long-range correlations and non-zero elliptic flow coefficients, raising questions about the minimal conditions required for QGP formation. However, one signature remains notably absent in small systems: jet quenching. Despite observing various collective phenomena, experiments have not identified energy loss of high-$p_{T}$ partons in pPb collisions Acharya and others (2018); Aad and others (2023). To further elucidate the system-size dependence of QGP signatures and to search for the possible onset of jet quenching in intermediate-sized systems, oxygen-oxygen (OO) and neon-neon (NeNe) collisions at a center-of-mass energy ($\sqrt{s_{\rm NN}}$) of 5.36 TeV per nucleon pair provided by the LHC in July, 2025 are of great interest Citron and others (2019); Brewer et al. (2021). These collision systems provide a crucial bridge between small and large systems, allowing precise control measurements of observables such as jet quenching and elliptic flow Lim et al. (2019); Sievert and Noronha-Hostler (2019); Huang et al. (2020); Huss et al. (2021); Jia and others (2024). Specifically, the role of $\alpha$ clusters in small nuclei including 16O has long been discussed and may be imprinted in the final-state flow measurements Rybczyński et al. (2018); Rybczyński and Broniowski (2019); Li et al. (2020); Summerfield et al. (2021); Ding et al. (2023); Wang et al. (2024); Zhang et al. (2025). Recently, comparing $v_{2}$ in NeNe to $v_{2}$ in OO collisions was highlighted to provide a precision test of the role of initial geometry and the QGP paradigm Giacalone and others (2025). Additionally, proton–oxygen (pO) collisions at a $\sqrt{s_{\rm NN}}=9.62$ TeV were provided, primarily to improve hadronic interaction models used in the interpretation of extensive air showers in cosmic-ray physicsDembinski et al. (2020), but they also present a unique opportunity to investigate QGP-related phenomena.
@@
Interpretation of high-energy nucleus collision data typically depends on the use of classical Glauber Monte Carlo (MCG) models to describe the initial condition Miller et al. (2007); d’Enterria and Loizides (2021). The general approach is to deduce properties of QCD matter by measuring the distributions of a variety of observables in AA collisions, and comparing them to the same measurements in pp collisions, where a QGP is not expected to be produced. To compare colliding systems of different sizes requires hence appropriate normalization of the measured distributions by using e.g. the number of participating nucleons ($N_{\rm part}$), the number of independent binary nucleon-nucleon (NN) collisions ($N_{\rm coll}$), the medium transverse area ($A_{\perp}$), or the overlap eccentricity $\epsilon_{n}$ (given by the $n^{\rm th}$ moments of its azimuthal spatial distribution).
@@
Widely used and publicly available MCG models are the TGlauberMC Loizides et al. (2015); Loizides (2016b); Loizides et al. (2018), which developed from the original PHOBOS implementation Alver et al. (2008) and GLISSANDO codes Broniowski et al. (2009); Rybczynski et al. (2014); Bożek et al. (2019). The Glauber initial condition is widely used in simulations of heavy-ion collisions as provided by TRENTO Moreland et al. (2015), Trajectum Nijs and van der Schee (2022b) and MUSIC Paquet et al. (2016), as well as by event generators such as HIJING Wang and Gyulassy (1991), AMPT Lin et al. (2005), URQMD Bass and others (1998) and EPOS Werner et al. (2010). An alternative description of the initial condition is provided by IP-Glasma Schenke et al. (2019).
@@
In this article, an updated version (v3.3) of the TGlauberMC model is used to predict properties of OO, NeNe, as well as pPb and PbPb collisions at 5.36 TeV, and pO collisions at 9.62 TeV. In particular, I focus on the calculation of the expected cross sections and discuss aforementioned associated initial-state quantities, such as $N_{\rm part}$, $N_{\rm coll}$ and eccentricity distributions for OO and NeNe collisions including the ratio of eccentricities for NeNe to OO as a proxy for the respective ratio of $v_{2}$ in these two systems. Finally, particle production at 5.36 TeV is discussed by outlining some features derived from the multiple parton interaction (MPI) picture Sjöstrand (2018), and use existing data at 5.36 TeV to predict the multiplicity in OO and NeNe collisions. As for previous versions of the model, the source code for TGlauberMC (v3.3) has been made publicly available at HepForge 92.
@@ II The Monte Carlo Glauber approach
Monte Carlo Glauber (MCG) calculations evaluate the cross section of two nuclei with $A$ and $B$ nucleons stochastically by distributing them in coordinate space according to the corresponding nuclear densities, separated by an impact parameter $b$ sampled from $\mathrm{d}\sigma/\mathrm{d}b\propto b$. For simplicity, the impact-parameter variable $b$ or $b_{\rm NN}$ denotes the distance between nuclei (or nucleons) in the transverse plane. The nuclear transverse profiles of the nuclei are usually approximated with parametrizations of charge density distributions extracted from low-energy electron-nucleus scattering experiments De Vries et al. (1987). For large nuclei, they are usually described by a 3-parameter Fermi (3pF) distribution
@@
with half-density radius $R$ and diffusivity $a$, and $w\neq 0$ describing small deviations from a spherical shape. Following the eikonal approximation Capel (2020), the collision is treated as a sequence of independent binary nucleon-nucleon collisions, i.e. the nucleons travel on straight-line trajectories and their interaction probability does not depend on the number of collisions they have suffered before. In the so-called “hard-sphere” (HS) approximation, an interaction takes place between two nucleons if the transverse distance $d$ (or generally the transverse nucleon–nucleon impact parameter $b_{\rm NN}$) between their centers satisfies $d<\sqrt{\sigma_{\rm NN}/\pi}$.
In the MCG approach, the number of participants and collisions are computed by counting the number of nucleons with at least one collision, and the total number of nucleon–nucleon collisions, respectively. The collision cross section is the fraction of simulated collisions with $N_{\rm coll}>0$ of all simulated collisions. Quantities such as the eccentricity Alver and others (2007), the triangularity Alver and Roland (2010), and higher moments Teaney and Yan (2011) of the collision region at impact parameter $b$ are given by
@@
where $n$ denotes the moment, $r=\sqrt{x^{2}+y^{2}}$ and $\psi=\tan^{-1}\frac{y}{x}$. The averages are computed by considering the central positions of either, participant nucleons or binary nucleon–nucleon collisions, or of an admixture of the two, and additionally may be smeared to model matter production Alver and others (2008).
@@ III Transverse overlap function
Several extensions to the hard-sphere approximation have been explored to account for the fact that the nucleon matter profile is not a black disc d’Enterria et al. (2010). A commonly applied extension, referred to as TRENTO Moreland et al. (2015), is to assume the nuclear matter profile of a nucleon to be a Gaussian with a certain width $w$, so that the nuclear overlap function
@@
is also Gaussian, i.e. $m=2$ and $B=2w$. For some settings (“bProfile = 3” Sjöstrand (2018)) PYTHIA assumes $B=1$ fm as characteristic size of the proton, and uses $m$ to interpolate between Exponential ($m\to 1$) and Gaussian ($m\to 2$) shape. The widely-used “Monash” tune sets $m=1.85$ Skands et al. (2014). HIJING Wang and Gyulassy (1991) uses an overlap function based on modified Bessel functions of the 3rd kind,
@@
with $\mu={1.5}/{\sqrt{\sigma_{\rm NN}}}$. In the Eikonal formalism, the probability to interact is then given as $p(b_{\rm NN})=1-\exp\left(-k\,T_{\rm NN}(b_{\rm NN})\right)$ where $k$ is obtained by requiring $2\pi\,\int p(b_{\rm NN})\,b_{\rm NN}\,{\rm d}b_{\rm NN}=\sigma_{\rm NN}$. At LHC collision energies, a practical parametrization Rybczyński and Wlodarczyk (2013) based on the Euler $\Gamma(z)$ and incomplete $\Gamma(\alpha,z)$ functions, in the following also called $\Gamma$-parametrization, is
@@
Figure 2: Nuclear charge density for 16O versus radius for various descriptions compared to data from electron scattering Sick and Mccarthy (1970). As explained in the text, “Oho2” is a parameterization based on the HO model, whose parameters are fit to the data, while “Odat” was already provided in Sick and Mccarthy (1970). “Osat” is the exact result of the NNLOsat calculation Somà et al. (2020)), and “Opar2” the respective 3pF fit. “Opar” and “Oho” profiles were previously included in TGlauberMC. The distributions are normalized as $\int\rho\,r^{2}\,{\rm d}r=Ze$. The respective $r_{\rm rms}$ values are given in the legend, denoted as $s$.
@@ IV Nuclear density profiles for oxygen and neon
The standard method to construct the nucleon configurations for MCG calculations is to sample the nucleon positions from $r^{2}\rho$, where the nuclear density $\rho$ may include angular deformations Loizides et al. (2015). To mimic a hard-core repulsion potential between nucleons, a minimum inter-nucleon separation ($d_{\rm min}$) of $0.4$ fm between their centers is enforced when sampling the positions of the nucleons inside a nucleus. In order to ensure that the center-of-mass of each constructed nucleus is at $(0,0,0)$, the nucleons are individually “shifted” by the necessary amount, after all nucleon positions were sampled. Both, imposing the hard-core repulsion as well as the re-centering, induce a bias from the original $\rho$ distribution. As discussed in Ref. Loizides et al. (2018), this can be overcome by a “reweighting” procedure, modifying the original $\rho$ distribution to counteract the bias. However, for $d_{\rm min}=0.4$ fm, and recentering via rotation around the $z$-axis (recentering option 4 in TGlauberMC), residual differences are negligible.
Table 1: Nuclear parameters for oxygen and neon parameterizations provided by TGlauberMC (v3.3) with given names as in the code. The 3pF parameterization is given in Eq. (1). The HO parameterization is given in Eq. (6). The last two configurations used deformed profiles, with deformation values as given in the comment field, and defined in Trajectum Giacalone and others (2025). The prefix “TR” indicates the profile is taken from Trajectum, where the name following the prefix denotes the system and version number as used in Trajectum. Included as “Osat” is also the exact profile from the ${\rm NNLO}_{\rm sat}$ chiral Hamiltonian calculation Somà et al. (2020). For completeness, “O” is also listed which uses explicit nucleon configurations Lonardoni et al. (2018) as introduced in Lim et al. (2019).
@@
Odat 16 8 2.608 0.513 -0.051 Original parametrization (3pF plus corrections) from Sick and Mccarthy (1970)
Opar 16 8 2.608 0.513 -0.051 3pF parameterization De Vries et al. (1987) (same as TR_OXYGEN_V3)
Opar2 16 8 1.850 0.497 0.912 3pF parameterization (TR_OXYGEN_V4, fit to ${\rm NNLO}_{\rm sat}$ Somà et al. (2020))
Oho 16 8 1.544 1.833 - Harmonic Oscillator De Vries et al. (1987)
Oho2 16 8 1.506 1.819 - Harmonic Oscillator (new fit to data from Sick and Mccarthy (1970))
Osat 16 8 - - - Exact NNLOsat profile from Somà et al. (2020), use instead of Opar2
O 16 8 - - - Nucleon configurations Lim et al. (2019) (same as in TGlauberMC v2.7)
Ne 20 10 2.805 0.571 - 2pF parametrization De Vries et al. (1987)
Ne2 20 10 2.740 0.572 - 2pF parametrization De Vries et al. (1987)
Ne3 20 10 2.791 0.698 -0.168 3pF parametrization De Vries et al. (1987)
NeTr2 20 10 2.805 0.571 - TR_NEON_V2, (same as Ne with $\beta_{2}=0.721$ from Pritychenko et al. (2016))
@@
Figure 2 shows the nuclear charge density for 16O for data from electron scattering Sick and Mccarthy (1970), including the original parameterization (“Odat”), as well as 3pF distributions with various parameters available in the literature (“Opar”, and “Opar2”). Ref. De Vries et al. (1987) also gives a parameterization based on the (modified) Harmonic Oscillator (HO) model
@@
labeled “Oho” in the figure. For this work, I also provide a new HO fit to the data, called “Oho2”, leading to $\alpha=1.506\pm 0.023$ and $1.819\pm 0.004$ with $\chi^{2}/n_{\rm dof}\approx 1$ assuming about 1% relative error on the data points. While it overall has a worse $\chi^{2}$ than “Odat”, it only has two parameters and can be used as alternative and allows systematic variation using the uncertainties provided on the fit parameters. To reflect the charge density, the distributions shown in Fig. 2 are normalized as $\int\rho\,r^{2}\,{\rm d}r=Ze$. A detailed list of all parameterizations shown in Fig. 2 is given in Tab. 1. The $r_{\rm rms}=\sqrt{\langle r\rangle}$ values are reported in the legend of the figure, and differ by nearly up to 5%. The reason is that the data do not constrain the tail of the distribution, which was found to affect the $r_{\rm rms}$ by a few percent, and thus experimental results between $2.65\pm 0.04$ and $2.73\pm 0.03$ have been reported Sick and Mccarthy (1970). Furthermore small differences can also be expected from inclusion of neutrons in the density distributions, and the use of point-like protons ignoring the proton form factor. However, the distribution “Opar2” which results from a fit to the ${\rm NNLO}_{\rm sat}$ calculation Somà et al. (2020) exhibits a significantly larger $r_{rms}$ than the data (and than reported in Somà et al. (2020)), and hence is not further considered here. Instead, the exact nuclear density obtained from the ${\rm NNLO}_{\rm sat}$ chiral Hamiltonian calculation Somà et al. (2020) is included and also shown in the figure. The nuclear charge density directly calculated by ${\rm NNLO}_{\rm sat}$ perfectly agrees with the data (also shown in the figure).
Table 2: Oxygen-16 and Neon-20 NLEFT and PGCM calculations taken from Trajectum Giacalone and others (2025).
@@
TR_OXYGEN_V14 PGCM, constraints imposed on projected states, without explicit $\alpha$ clustering ($C=0$)
TR_OXYGEN_V15 PGCM, constraints imposed on projected states, including explicit $\alpha$ clustering ($C=1$)